OpenStudy (anonymous):

Which shows the mixed numbers in order from greatest to least? four and five-eighths, four and five-sixths, four and three-fourths Choose one answer. a. four and five-sixths, four and three-fourths, four and five-eighths b. four and three-fourths, four and five-sixths, four and five-eighths c. four and five-eighths, four and three-fourths, four and five-sixths d. four and five-sixths, four and five-eighths, four and three-fourths

OpenStudy (anonymous):

The only one that makes sense is c

OpenStudy (anonymous):

Because four and five-eighths is the greatest

OpenStudy (anonymous):

The easiest way to figure it out is to ignore the 4 because it does not effect the fractions.

OpenStudy (anonymous):

if you convert them all to x/48 then you can figure it out fairly easily

OpenStudy (anonymous):

so 8*6=48 and 5*6=30 so 5/8=30/48

OpenStudy (anonymous):

8*6=48 and 5*8=40 so 5/6=40/48

OpenStudy (anonymous):

4*12=48 and 3*12=36 so 3/4=36/48

OpenStudy (gabylovesyou):

C is not correct since as he demonstrates, 5/6 is greater than 5/8 :] try again

OpenStudy (anonymous):

So it would be A. Right?

OpenStudy (anonymous):

which makes it 40/48, 36/48, 30/48, or 5/6, 3/4, 5/8 so it is A, yes

OpenStudy (anonymous):

Thank you :D

OpenStudy (anonymous):

No problem :)